OpenStudy (anonymous):

How to solve this equation using LCD method? Step by step explanation much appreciated

OpenStudy (anonymous):

\[\frac{ 3 }{ x+2 } +\frac{ 1 }{ 2x } = \frac{ 4 }{ x+2 }\]

OpenStudy (usukidoll):

LCD is (x+2)(2x) multiply the numerator and denominator with that. Then cancel out whatever is in common. for example on the first fraction 2x is missing so multiply the 2x on the numerator and denominator second part is (x+2) . hmm you should subtract the 4/x+2 and combine like terms before doing lcd

OpenStudy (anonymous):

Could you explain how to solve the equation completely?

OpenStudy (anonymous):

3/(x+2) + 1/2x = 4/(x+2) so, to find the Least Common Denominator, in this case just multiply (x+2) x 2x. you always have to multiply the numerator too when multiplying the denominator. You just can't leave out the numerator. so, [3(2x)/(x+2)(2x)] + [1(x+2)/(x+2)(2x)] = 4(2x)/(x+2)(2x)

OpenStudy (anonymous):

Would you distribute and get: 6x/(x+2)(2x) + x+2/(x+2)(2x) = 8x/(x+2)(2x) Is this right? and if it is what do I do next?

OpenStudy (anonymous):

I solved the problem! Thank you for all the help ^-^